OpenStudy (darkbluechocobo):

Use substitution to solve the system of equations. Enter your answer as an ordered pair; example: (4, -5) y = 3x + 5 x = -3y - 5

OpenStudy (ranga):

Take y from the first equation substitute it in the second equation and solve for x. Then put x value in one of the equations and find y.

OpenStudy (anonymous):

x=-3(-5)+5 x=-15+5 x=-10 I think that's half of the steps right there.

OpenStudy (darkbluechocobo):

ooo i see

OpenStudy (darkbluechocobo):

so then you would put -10 into the y solution

OpenStudy (anonymous):

Yep I think so

OpenStudy (darkbluechocobo):

so y=3(-10)+5

OpenStudy (darkbluechocobo):

y=-25

OpenStudy (darkbluechocobo):

actually I just realized. That was an example, it is not actually used in the problem

OpenStudy (darkbluechocobo):

the ordered pair (4,5) is not used so what would you do now??

OpenStudy (anonymous):

I think you are supposed to use x which is 4 in the problem to? I can't remember but. Yeah just plug -10 in and I think you are good!

OpenStudy (anonymous):

Then you find x and y

OpenStudy (darkbluechocobo):

I believe you would do something like x= -3(3x+5)-5

OpenStudy (***[isuru]***):

nope.... Enter your answer as an ordered pair; example: (4, -5) the ordered pair is only an example to how to present ur answer... substitution is y = 3x + 5 x = -3y - 5 x = -3(3x + 5) -5 then solve it

OpenStudy (anonymous):

Ah ok sorry about that.

OpenStudy (darkbluechocobo):

Yes I thought it was that :p

OpenStudy (darkbluechocobo):

so x=-9x-20?

OpenStudy (***[isuru]***):

yep!

OpenStudy (***[isuru]***):

nway... u got the correct idea... that's the important part...

OpenStudy (darkbluechocobo):

so then how would you put it into an ordered pair solution o.o

OpenStudy (ranga):

You solve for x first. Then find y. Put it as (x, y) and that is how you put it as an ordered pair. What are you getting for x?

OpenStudy (***[isuru]***):

ordered pair ---- (x value , y value )

OpenStudy (darkbluechocobo):

Yes but how do you just seclude it to one number if this is what you get for x x=-9x-20?

OpenStudy (***[isuru]***):

solve it for x

OpenStudy (ranga):

Oh, you have to solve for x x = -9x - 20 Add 9x to both sides. What do you get?

OpenStudy (***[isuru]***):

x + 9x = -20 10x = -20 x = ?

OpenStudy (darkbluechocobo):

x=2

OpenStudy (darkbluechocobo):

well that was a facepalm moment *facepalm*

OpenStudy (ranga):

watch out for the sign

OpenStudy (***[isuru]***):

u have messed up with -,+signs

OpenStudy (darkbluechocobo):

-2

OpenStudy (***[isuru]***):

yep

OpenStudy (darkbluechocobo):

so y=3(-2)+5

OpenStudy (***[isuru]***):

go on...

OpenStudy (ranga):

OK. Put x = -2 in any one of the equations and find y. What do you get for y?

OpenStudy (darkbluechocobo):

-6+5 is -1 y=-1

OpenStudy (***[isuru]***):

yep

OpenStudy (ranga):

Now put it as an ordered pair

OpenStudy (darkbluechocobo):

so (-2,-1)

OpenStudy (***[isuru]***):

ya got it!!

OpenStudy (ranga):

Yes!

OpenStudy (darkbluechocobo):

:D

OpenStudy (darkbluechocobo):

Thank you guys

OpenStudy (darkbluechocobo):

If I could give more than one best response >.<

OpenStudy (***[isuru]***):

u r welcome !!

OpenStudy (***[isuru]***):

ranga --> darkblue me ---> ranga best response

OpenStudy (darkbluechocobo):

Could You help me with some more. Because like this is all about system of equations and its not my best point of algebra

OpenStudy (***[isuru]***):

ok... sure if i can

OpenStudy (darkbluechocobo):

Alright Thank you >.< I just have issues in this part of algebra

OpenStudy (darkbluechocobo):

Classify the system and determine the number of solutions. (Check all that apply.) $x - 2y = -8$ $4x = 8y - 56$ inconsistent consistent no solution infinitely many solutions

OpenStudy (darkbluechocobo):

ignore the dollar signs haha

OpenStudy (darkbluechocobo):

x=2y-14

OpenStudy (***[isuru]***):

is there any specific way to solve this ....like using substitution... or can we solve it by any method?

OpenStudy (darkbluechocobo):

By any method I believe saying it doesn't say a specific type

OpenStudy (***[isuru]***):

x=2y-14 -----(1) x - 2y = -8 x = 2y - 8 -----(2) so.... no solutions... (1) = (2) 2y -14 = 2y -8 y will cancel out.. u can't solve this one...

OpenStudy (darkbluechocobo):

so no solution and inconsistent

OpenStudy (***[isuru]***):

yep

OpenStudy (darkbluechocobo):

Thank you again :3 Use substitution to solve the system of equations. Enter your answer as an ordered pair; example: (4, -5) 5x + 6y = -9 2x - 2 = -y With this one can you just correct me if I get it wrong :p

OpenStudy (darkbluechocobo):

y=-2x+2

OpenStudy (***[isuru]***):

correct

OpenStudy (darkbluechocobo):

5x=6y+9 x=1.2y+1.8

OpenStudy (darkbluechocobo):

or do you put y in before you divide it by 5

OpenStudy (***[isuru]***):

use substitution

OpenStudy (***[isuru]***):

5x = 6(-2x+2) + 9

OpenStudy (darkbluechocobo):

-12x+12+9

OpenStudy (darkbluechocobo):

-2.4x+2.4+1.8

OpenStudy (darkbluechocobo):

x=-2.4x+4.2

OpenStudy (darkbluechocobo):

3.4x=4.2

OpenStudy (darkbluechocobo):

-3.4*

OpenStudy (darkbluechocobo):

I get a repeating number

OpenStudy (***[isuru]***):

5x + 6y = -9 5x + 6(-2x+2) = -9 5x - 12x + 12 = -9 -7x = -21 x = 3 u have miss the step here 5x + 6y = -9 ----- > 5x = 6y - 9 ; it should be 5x = -6y - 9

OpenStudy (darkbluechocobo):

so 12x-12+9

OpenStudy (darkbluechocobo):

12x-3

OpenStudy (***[isuru]***):

12x -12-9

OpenStudy (***[isuru]***):

12x - 21

OpenStudy (darkbluechocobo):

*facepalm*

OpenStudy (***[isuru]***):

lol

OpenStudy (darkbluechocobo):

7x=-21

OpenStudy (darkbluechocobo):

x=-3

OpenStudy (***[isuru]***):

nope

OpenStudy (***[isuru]***):

5x = 12x - 21 5x - 12 x = -21 -7x = -21

OpenStudy (darkbluechocobo):

UHG *flips table* CURSE YOU NEGATIVES

OpenStudy (darkbluechocobo):

so x= 3

OpenStudy (***[isuru]***):

yeah!!

OpenStudy (darkbluechocobo):

2(3)-2=y

OpenStudy (darkbluechocobo):

6-2=4 y=4

OpenStudy (darkbluechocobo):

(3,4)

OpenStudy (***[isuru]***):

y = -2x + 2 = -2(3) + 2 = -6 + 2 = -4

OpenStudy (darkbluechocobo):

._.

OpenStudy (darkbluechocobo):

Its official my mind hates me

OpenStudy (darkbluechocobo):

(3,-4)

OpenStudy (***[isuru]***):

yes!!!

OpenStudy (darkbluechocobo):

Haha Thank you again and sorry for my negative troubles >.<

OpenStudy (darkbluechocobo):

Classify the system and determine the number of solutions. (Check all that apply.) $8x + 12y = 60$ $2x + 3y = -24$ no solution consistent inconsistent infinitely many solutions

OpenStudy (darkbluechocobo):

again sorry about the dollar signs ._.

OpenStudy (***[isuru]***):

no problem... but watch out for negatives...

OpenStudy (***[isuru]***):

2x = -24 - 3y 8x + 12y = 60 4(2x) + 12y = 60 4(-24 - 3y) + 12y = 60 -96 -12y + 12 y = 60 so... it's the same as before....

OpenStudy (darkbluechocobo):

Now substitution ._. Use elimination to solve the system of equations. Enter your answer as an ordered pair; example: (4, -5) 1/2x+y=4 -2x-2y=-6

OpenStudy (darkbluechocobo):

Also I really appreciate your help :3 Thank you